OpenStudy (mallorysipp234):

really confused. Given the function f(x) = 0.5(3)x, what is the value of f−1(7)? 0.565 1.140 1.771 2.402

OpenStudy (mallorysipp234):

the -1 is an exponent to the F like so: \[f^-1 \]

OpenStudy (akashdeepdeb):

Is that f(x) = \(0.5*3x\) Or is it f(x) = \(0.5^3x\) ?

OpenStudy (mallorysipp234):

wait, its 0.5 x (3) ^ x

OpenStudy (akashdeepdeb):

\(f^{-1}(x)\) means that it is an inverse of the function \(f(x)\) So, When \(f(x) = 0.5 * 3 ^ x\) we can write that as \(y = 0.5 * 3 ^ x\) Then the inverse of this would be [by switching the x and the y values] \(x = 0.5 * 3 ^ y\) So \(\frac{x}{0.5} = 3 ^ y\) Or \[y =\log_{3} \frac{x}{0.5} \] Getting this? :)

OpenStudy (mallorysipp234):

yeah i understand that.

OpenStudy (akashdeepdeb):

Now just substitute the value of 7 in place of x and then find y = log_3 (14) Got it right? :D

OpenStudy (mallorysipp234):

my calculator is telling me that log_3 (14) is 6.679 but that's none of my answers. im still confused. :(

OpenStudy (akashdeepdeb):

Try again, my prediction is the answer should be greater than 2 and lesser than 3. 3 ^ 6.679 = 729 ++

OpenStudy (akashdeepdeb):

Do you see why? :D To check you can just take random natural numbers and find a suitable range. For eg. we want 3 ^ (something) = 54 If that something is 2 then 3^3 = 27 And if it is 4 then 3 ^ 4 = 81 Then we can definitively say that that something lies between (3 --> 4) Getting me? :)

OpenStudy (mallorysipp234):

yes

OpenStudy (akashdeepdeb):

So the answer of this question should be?

OpenStudy (mallorysipp234):

2.402 would be the only thing that is in between the range of greater than two lesser than three.

OpenStudy (akashdeepdeb):

Excellent, you got it! (y)

OpenStudy (mallorysipp234):

can you help with one more? similar to this one.

OpenStudy (akashdeepdeb):

I'd suggest you try it first and if you have any question after that, then ask! :D

OpenStudy (mallorysipp234):

this one has to do with a log.

OpenStudy (akashdeepdeb):

Fine, let me see if i can help. :)

OpenStudy (mallorysipp234):

If f(x) = log2 (x + 4), what is f−1(3)? 0 2 4 8

OpenStudy (mallorysipp234):

@AkashdeepDeb

OpenStudy (akashdeepdeb):

Alright so what was the initial procedure? What did we do in the beginning?

OpenStudy (akashdeepdeb):

@mallorysipp234 ?

OpenStudy (mallorysipp234):

switch the x and the y.

OpenStudy (akashdeepdeb):

Yes, so what do you get after doing that?

OpenStudy (mallorysipp234):

wouldn't it be log_2 (x+4) = 3

OpenStudy (akashdeepdeb):

f(x) = log2 (x + 4) y = log2 (x+4) SWITH THEM NOW. :D x = log2 (y+4) So NOW we have to find y when x=3 3 = log2 (y+4) [And thus we can find y] Now do you know how to convert this function to exponential form [a ^ b = c form]

OpenStudy (mallorysipp234):

no :(

OpenStudy (akashdeepdeb):

Hint: \(a^b = c \) can be written as \(log_{a}~ c = b\) try now? :D

OpenStudy (akashdeepdeb):

Learn that ^ It is very important in Math. :)

OpenStudy (mallorysipp234):

so log_2(3) = (y+4) ?????

OpenStudy (mallorysipp234):

alright ill write it down in my notes

OpenStudy (akashdeepdeb):

Not quite. 3 = log2 (y+4) in a better format is \( log_2 (y+4) = 3\) So when 2 is the base and 3 is the power. We can say that \(2^3 = 8 = (y+4)\) Did you get this? If you didn't, ask, I'll explain it again.

OpenStudy (mallorysipp234):

yeah! that makes sense.

OpenStudy (mallorysipp234):

so y is 4.

OpenStudy (akashdeepdeb):

Correct.

OpenStudy (mallorysipp234):

alright! (: i get it.. thank you!

OpenStudy (akashdeepdeb):

:)